# Continuum limit of the mobility edge and taste-degeneracy effects in high-temperature lattice QCD with staggered quarks

### 05 Dec 2023 v1

@@ Abstract
We study the effects of taste degeneracy on the continuum scaling of the localization properties of the staggered Dirac operator in high-temperature QCD using numerical simulations on the lattice, focusing in particular on the position of the mobility edge separating localized and delocalized modes at the low end of the spectrum. We find that, if the continuum limit is approached at fixed spatial volume, the restoration of taste symmetry leads to sizeable systematic effects on estimates for the mobility edge obtained from spectral statistics, which become larger and larger as the lattice spacing is decreased. Such systematics, however, are found to decrease if the volume is increased at fixed lattice spacing. Using an independent analysis based directly on the properties of the Dirac eigenvectors, that are unaffected by taste degeneracy, we show that taking the thermodynamic limit before the continuum limit leads to the correct result for the mobility edge estimated from spectral statistics. This allows us to obtain a controlled continuum extrapolation of the mobility edge for the first time.
@@ 1 Introduction
On the one hand, a lattice Dirac operator in the background of the fluctuating gauge fields appearing in the path-integral formulation of gauge theories is formally a (sparse) random matrix. As such, its eigenvalues are expected to display the universal types of correlations appearing in these systems [37, 38]. This allows one to identify the mobility edge as the point in the spectrum where the spectral statistics switches between the universal type corresponding to localized modes, and the universal type corresponding to delocalized modes (which depends only on the symmetry class of the lattice Dirac operator according to the random matrix theory classification [38, 39]). On the other hand, in four space-time dimensions the staggered operator describes four “tastes” of fermions that become exactly degenerate in the continuum limit, which in turn leads the staggered spectrum to develop degenerate quartets of eigenvalues as the lattice spacing decreases towards zero. The obvious additional correlations between the nearly-degenerate eigenvalues appearing on fine lattices then spoil the expected universal behavior. These deviations are difficult to control theoretically, and make it more difficult to reliably determine the localization properties of eigenmodes using spectral statistics on finer lattices. This issue has been known for a long time [40, 41, 4], [^1] For certain gauge groups and choices of representation for the fermions, the symmetries of the continuum Dirac operator differ from those of the staggered one [38]: In these cases, the deviation from the expected universal behavior is actually needed, in order for the statistical properties of the lattice spectrum to approach those of the continuum spectrum (once the degeneracy is removed). However, deviations from the universal behavior will appear even when the staggered and continuum operator are in the same symmetry class. but it has never been fully addressed in the lattice literature so far. A careful lattice study would be extremely interesting to assess possible systematics affecting the calculation of the mobility edge.
@@
The discussion above shows that if we take the thermodynamic limit before the continuum one, we can expect the distortions of the spectral statistics due to taste degeneracy to disappear, and the expected universal behavior based on the exact symmetries of the staggered operator should emerge. This would in turn allow one to employ standard methods to determine the localization properties of the eigenmodes using spectral statistics, that can then be extrapolated to the continuum. On the other hand, since the residual fermion doubling of sea staggered quarks is usually dealt with using the so-called “rooting trick” [50, 51, 52], one expects the correct order of limits to be continuum first and thermodynamic after. In this way, one recovers an exact taste degeneracy in the sea lattice action, so that taking the fourth root of the determinant in the path integral is justified. However, as explained above, the uncontrolled effects of taste degeneracy on the spectral statistics of the valence staggered operator make the determination of localization properties and of the position of the mobility edge unreliable on finer lattices if the volume is not large enough. In this context, a non-trivial question is whether one can actually exchange the order of limits and still obtain the correct results.
@@
In this paper we study in detail the effects of taste degeneracy on the statistical properties of the staggered Dirac spectrum, and how they affect the determination of the mobility edge using spectral statistics, with the ultimate goal of providing a controlled investigation of the behavior of the mobility edge in the continuum limit. After briefly reviewing the properties of the staggered operator, and discussing localization and how to detect it using both eigenvector properties and the statistical properties of the spectrum, in Sec. 2 we discuss the complications due to the formation of nearly degenerate eigenvalue multiplets, and how the lattice spacing and the lattice volume affect them. In Sec. 3 we present numerical determinations of the mobility edge in the staggered spectrum of high-temperature QCD for various lattice spacings and volumes, obtained by a standard analysis of spectral statistics, and study the dependence on the lattice spacing and on the lattice volume. We then compare the results with a determination based on a direct study of eigenvector properties, unaffected by taste degeneracy. Finally, we discuss the extrapolation to the continuum limit. We draw our conclusions in Sec. 4.
@@ 2.1 The staggered lattice Dirac operator
where $U_{\mu}(x)\in\mathrm{SU}(3)$, with $x=(x_{1},x_{2},x_{3},x_{4})$ and $\mu=1,\ldots,4$, is a gauge link variable attached to the link connecting the sites $x$ and $x+a\hat{\mu}$ of a $N_{s}\times N_{s}\times N_{s}\times N_{t}$ hypercubic lattice with lattice spacing $a$, $\mathrm{T}_{\mu}$ are unit translation operators with periodic boundary conditions in the spatial directions $\mu=1,2,3$ and antiperiodic boundary conditions in the temporal direction $\mu=4$, and $\eta_{\mu}$ are the staggered phases, $\eta_{\mu}(x)=(-1)^{\sum_{\alpha<\mu}x_{\alpha}}$. The operator $D^{\mathrm{stag}}$ is anti-Hermitean, so its eigenmodes obey
@@ 2.3 Spectral statistics
The discussion above is valid only in the limit of infinite volume, where any overlap between different localized modes becomes entirely negligible, and where delocalized modes can fully spread out on a system of infinite size. In any finite volume there are instead corrections to the Poisson or RMT-type behavior, that tend to zero as the volume is increased. In the presence of a mobility edge one then expects $p(s;\lambda,N_{s})$ to interpolate between a near-exponential behavior and a near-RMT behavior, passing through the critical behavior at the mobility edge, where the properties of the eigenmodes are expected to be scale-invariant. Then, as $N_{s}$ is increased, $p\to p_{\mathrm{P}}$ in the localized spectral region, $p\to p_{\mathrm{RMT}}$ in the delocalized spectral region, and $p=p_{*}$ at the mobility edge $\lambda=\lambda_{c}$. This can be used to determine $\lambda_{c}$ and the critical exponents of the Anderson transition by means of a finite size scaling analysis [63], as was done in Ref. [5] for QCD. Alternatively, if the critical distribution $p_{*}$ is known (as is the case, to same extent, for the unitary class [64]), one can use it to determine $\lambda_{c}$ as the point where the spectral statistics matches the critical one. This can be done very efficiently using a single lattice volume: by virtue of the scale-invariance of the critical point, such an estimate is expected to suffer only from very little finite-size effects.
@@ 2.4 Effects of taste degeneracy
The discussion above ignored entirely the possible presence of additional correlations between eigenvalues due to peculiarities in the structure of the random Hamiltonian. Such correlations can be induced by the presence of an approximate symmetry, leading to near-degenerate multiplets of eigenvalues. In this case, the corresponding symmetry-breaking parameter controls the typical splitting $\delta$ within the would-be-degenerate multiplets, which, in turn, gives the spectral scale at which their effects are felt. When this scale is smaller than or comparable to the expected distance $\Delta=1/\rho_{0}$ between neighboring eigenvalues, it affects the unfolded level spacing distribution, skewing it towards lower values: indeed, neighboring levels belonging to the same approximate multiplet prefer to stay closer than what one would expect simply based on the inverse of the spectral density, as their distance is controlled by a different mechanism with a corresponding smaller spectral scale. More precisely, one has clearly separated multiplets of $n$ eigenvalues if their size $(n-1)\delta$ is much smaller than the typical distance $n/\rho_{0}=n\Delta$ between them, i.e., if $\delta\ll\Delta$. In this case, out of $n$ level spacings, $n-1$ are of order $\delta$, and one is of order $n\Delta-(n-1)\delta$. This means that the average level spacing results from the averaging of fluctuations around two different typical values; for this reason, one expects also a larger variance for the unfolded level spacing distribution.
@@
The staggered operator provides precisely an example of the situation discussed above: while in the continuum limit it has an exact symmetry under the exchange of degenerate tastes, on a finite lattice this symmetry is only approximate. This manifests in the formation of nearly-degenerate eigenvalue multiplets as one gets sufficiently close to the continuum. The symmetry-breaking parameter is here the lattice spacing $a$, and the splittings $\delta$ in the nearly-degenerate eigenvalue multiplets are of order $\delta\sim(a\Lambda)^{2}\Lambda$ (see, e.g., Ref. [65]), with $\Lambda$ some physical mass scale. In the thermodynamic limit at fixed $a$, the argument above applies: the multiplet structure gets washed out, and its effects on short-range eigenvalue correlations disappear. On the other hand, as $a\to 0$ the splittings decrease and multiplets become more degenerate, and more distinguishable from each other: this effect competes with the wash-out effect of the thermodynamic limit, and so the order in which the continuum and the thermodynamic limits are taken becomes important.
@@
However, even if this is the case, this approach rapidly gets numerically very intensive both for the generation of gauge configurations and for the diagonalization of the staggered operator as one gets closer to the continuum, until it becomes impossible to reach sufficiently large aspect ratios to see the plateau, and the extrapolation of the results for the mobility edge towards the thermodynamic limit introduces larger systematic effects. As the lattice spacing is reduced the statement that the approach based on spectral statistics is numerically more efficient becomes questionable.

### 12 Mar 2024 v2

@@ Abstract
We study the effects of taste degeneracy on the continuum scaling of the localization properties of the staggered Dirac operator in high-temperature QCD using numerical simulations on the lattice, focusing in particular on the position of the mobility edge separating localized and delocalized modes at the low end of the spectrum. We find that, if the continuum limit is approached at fixed spatial volume, the restoration of taste symmetry leads to sizeable systematic effects on estimates for the mobility edge obtained from spectral statistics, which become larger and larger as the lattice spacing is decreased. Such systematics, however, are found to decrease if the volume is increased at fixed lattice spacing. We argue that spectral statistics estimate correctly the position of the mobility edge in the thermodynamic limit at fixed spacing, and support this with an independent numerical analysis based directly on the properties of the Dirac eigenvectors, that are unaffected by taste degeneracy. We then provide a theoretical argument justifying the exchange of the thermodynamic and continuum limits when studying localization. This allows us to use spectral statistics to determine the position of the mobility edge, and to obtain a controlled continuum extrapolation of the mobility edge for the first time.
@@ 1 Introduction
On the one hand, a lattice Dirac operator in the background of the fluctuating gauge fields appearing in the path-integral formulation of gauge theories is formally a (sparse) random matrix. As such, its eigenvalues are expected to display the universal types of correlations appearing in these systems [37, 38]. This allows one to identify the mobility edge as the point in the spectrum where the spectral statistics switch between the universal type corresponding to localized modes, and the universal type corresponding to delocalized modes (which depends only on the symmetry class of the lattice Dirac operator according to the random matrix theory classification [38, 39]). On the other hand, in four space-time dimensions the staggered operator describes four “tastes” of fermions that become exactly degenerate in the continuum limit, which in turn leads the staggered spectrum to develop degenerate quartets of eigenvalues as the lattice spacing decreases toward zero. The obvious additional correlations between the nearly-degenerate eigenvalues appearing on fine lattices then spoil the expected universal behavior. These deviations are difficult to control theoretically, and make it more difficult to reliably determine the localization properties of eigenmodes using spectral statistics on finer lattices. This issue has been known for a long time [40, 41, 4], [^1] For certain gauge groups and choices of representation for the fermions, the symmetries of the continuum Dirac operator differ from those of the staggered one [38]: In these cases, the deviation from the expected universal behavior is actually needed, in order for the statistical properties of the lattice spectrum to approach those of the continuum spectrum (once the degeneracy is removed). However, deviations from the universal behavior will appear even when the staggered and continuum operator are in the same symmetry class. but it has never been fully addressed in the lattice literature so far. A careful lattice study would be extremely interesting to assess possible systematics affecting the calculation of the mobility edge.
@@
The discussion above shows that if we take the thermodynamic limit before the continuum one, we can expect the distortions of the spectral statistics due to taste degeneracy to disappear, and the expected universal behavior based on the exact symmetries of the staggered operator should emerge. This would in turn allow one to employ standard methods to determine the localization properties of the eigenmodes using spectral statistics, that can then be extrapolated to the continuum. On the other hand, since the residual fermion doubling of sea staggered quarks is usually dealt with using the so-called “rooting trick” [50, 51, 52], one expects the correct order of limits to be continuum first and thermodynamic afterward. In this way, one recovers an exact taste degeneracy in the sea lattice action, so that taking the fourth root of the determinant in the path integral is justified. However, as explained above, the uncontrolled effects of taste degeneracy on the spectral statistics of the valence staggered operator make the determination of localization properties and of the position of the mobility edge using these quantities unreliable on finer lattices if the volume is not large enough. In this context, a non-trivial question is whether one can actually exchange the order of limits and still obtain the correct results. It is important to note that exchanging limits could be allowed when determining the mobility edge even though it is not when studying spectral statistics. In fact, the mobility edge depends only on the localization properties of the eigenmodes, which are well defined entirely independently of the statistical properties of the spectrum. While localization properties and spectral statistics are related, it is the former that strongly affect the latter, and not the other way around.
@@
In this paper we study in detail the effects of taste degeneracy on the statistical properties of the staggered Dirac spectrum, and how they affect the determination of the mobility edge using spectral statistics, with the ultimate goal of providing a controlled investigation of the behavior of the mobility edge in the continuum limit. After briefly reviewing the properties of the staggered operator, and discussing localization and how to detect it using both eigenvector properties and the statistical properties of the spectrum, in Sec. 2 we discuss the complications due to the formation of nearly degenerate eigenvalue multiplets, and how the lattice spacing and the lattice volume affect them. We then argue that the continuum and thermodynamic limits can be exchanged when studying the mobility edge. In Sec. 3 we present numerical determinations of the mobility edge in the staggered spectrum of high-temperature QCD for various lattice spacings and volumes, obtained by a standard analysis of spectral statistics, and study the dependence on the lattice spacing and on the lattice volume. We then compare the results with a determination based on a direct study of eigenvector properties, unaffected by taste degeneracy, and with a heuristic estimate from the spectral statistics of a suitably “taste-symmetrized” spectrum, where the approximate taste degeneracy is removed by hand. Finally, we discuss the extrapolation to the continuum limit. We draw our conclusions in Sec. 4.
@@ 2.1 The staggered lattice Dirac operator
where $U_{\mu}(x)\in\mathrm{SU}(3)$, with $x=(x_{1},x_{2},x_{3},x_{4})$ and $\mu=1,\ldots,4$, is a gauge link variable attached to the link connecting the sites $x$ and $x+a\hat{\mu}$ of a $N_{s}\times N_{s}\times N_{s}\times N_{t}$ hypercubic lattice with lattice spacing $a$; $\mathrm{T}_{\mu}$ are unit translation operators with periodic boundary conditions in the spatial directions $\mu=1,2,3$ and antiperiodic boundary conditions in the temporal direction $\mu=4$; and $\eta_{\mu}$ are the staggered phases, $\eta_{\mu}(x)=(-1)^{\sum_{\alpha<\mu}x_{\alpha}}$. The operator $D^{\mathrm{stag}}$ is anti-Hermitean, so its eigenmodes obey
@@ 2.3 Spectral statistics
The discussion above is valid only in the limit of infinite volume, where any overlap between different localized modes becomes entirely negligible, and where delocalized modes can fully spread out on a system of infinite size. In any finite volume there are instead corrections to the Poisson or RMT-type behavior, that tend to zero as the volume is increased. In the presence of a mobility edge one then expects $p(s;\lambda,N_{s})$ to interpolate between a near-exponential behavior and a near-RMT behavior, passing through the critical behavior at the mobility edge, where the properties of the eigenmodes are expected to be scale-invariant. Then, as $N_{s}$ is increased, $p\to p_{\mathrm{P}}$ in the localized spectral region, $p\to p_{\mathrm{RMT}}$ in the delocalized spectral region, and $p=p_{*}$ at the mobility edge $\lambda=\lambda_{c}$. This can be used to determine $\lambda_{c}$ and the critical exponents of the Anderson transition by means of a finite size scaling analysis [63], as was done in Ref. [5] for QCD. Alternatively, if the critical distribution $p_{*}$ is known (as is the case, to some extent, for the unitary class [64]), one can use it to determine $\lambda_{c}$ as the point where the spectral statistics matches the critical one. This can be done very efficiently using a single lattice volume: by virtue of the scale-invariance of the critical point, such an estimate is expected to suffer only from very little finite-size effects.
@@ 2.4 Effects of taste degeneracy
The discussion above ignored entirely the possible presence of additional correlations between eigenvalues due to peculiarities in the structure of the random Hamiltonian. Such correlations can be induced by the presence of an approximate symmetry, leading to near-degenerate multiplets of eigenvalues. In this case, the corresponding symmetry-breaking parameter controls the typical splitting $\delta$ within the would-be-degenerate multiplets, which, in turn, gives the spectral scale at which their effects are felt. When this scale is smaller than or comparable to the expected distance $\Delta=1/\rho_{0}$ between neighboring eigenvalues, it affects the unfolded level spacing distribution, skewing it toward lower values: Indeed, neighboring levels belonging to the same approximate multiplet prefer to stay closer than what one would expect simply based on the inverse of the spectral density, as their distance is controlled by a different mechanism with a corresponding smaller spectral scale. More precisely, one has clearly separated multiplets of $n$ eigenvalues if their size $(n-1)\delta$ is much smaller than the typical distance $n/\rho_{0}=n\Delta$ between them, i.e., if $\delta\ll\Delta$. In this case, out of $n$ level spacings, $n-1$ are of order $\delta$, and one is of order $n\Delta-(n-1)\delta$. This means that the average level spacing results from the averaging of fluctuations around two different typical values; for this reason, one expects also a larger variance for the unfolded level spacing distribution.
@@
The staggered operator provides precisely an example of the situation discussed above: while in the continuum limit it has an exact symmetry under the exchange of degenerate tastes, on a finite lattice this symmetry is only approximate. This manifests in the formation of nearly-degenerate eigenvalue multiplets as one gets sufficiently close to the continuum. The symmetry-breaking parameter is here the lattice spacing $a$, and the splittings $\delta$ in the nearly-degenerate eigenvalue multiplets are of order $\delta\sim(a\Lambda)^{2}\Lambda$ (see, e.g., Ref. [65, 66]), with $\Lambda$ some physical mass scale. In the thermodynamic limit at fixed $a$, the argument above applies: the multiplet structure gets washed out, and its effects on short-range eigenvalue correlations disappear. On the other hand, as $a\to 0$ the splittings decrease and multiplets become more degenerate, and more distinguishable from each other: this effect competes with the wash-out effect of the thermodynamic limit, and so the order in which the continuum and the thermodynamic limits are taken becomes important.
@@
However, even if this is the case, this approach rapidly gets numerically very intensive both for the generation of gauge configurations and for the diagonalization of the staggered operator as one gets closer to the continuum, until it becomes impossible to reach sufficiently large aspect ratios to see the plateau, and the extrapolation of the results for the mobility edge toward the thermodynamic limit introduces larger systematic effects. As the lattice spacing is reduced the statement that the approach based on spectral statistics is numerically more efficient becomes questionable.